# Movie2Story: A framework for understanding videos and telling stories in the form of novel text

### v1

@@ 4.3 Audio
4.3.1 Speaker Recognition
Speaker recognition is an important branch of speech technology aimed at identifying or verifying the identity of a speaker based on their speech features. It extracts the unique identity information of the speaker by analyzing the acoustic features of the speech signal, and performs matching and recognition. Speaker recognition technology plays an important role in fields such as security monitoring, intelligent voice assistants, and identity authentication. In this article, we use distinguishing different speakers to achieve audio and video character matching for accurate text description.
1. Feature Extraction:
To capture the acoustic characteristics of each audio segment, we employ the pyannote.audio toolkit, which offers a range of feature extraction techniques. This includes standard methods like Mel-frequency cepstral coefficients (MFCCs) and spectrograms, providing a rich representation of the audio signal. Additionally, pyannote.audio supports on-the-fly data augmentation, enhancing the robustness and generalization of the models.
2. Speaker Embedding:
@@
The core of our speaker recognition system lies in the Speaker Embedding module, powered by pyannote.audio. This module leverages deep learning models to extract unique and discriminative features from each speaker’s voice. These speaker embeddings serve as compact representations that capture the essential aspects of a speaker’s identity, facilitating accurate recognition and matching.
pyannote.audio provides a unified framework for training sequence labeling models for tasks like Voice Activity Detection, Speaker Change Detection, and Overlapping Speech Detection. By employing recurrent neural networks and attention mechanisms, these models effectively learn to identify and segment different speakers within the audio stream.
To further improve speaker recognition performance, pyannote.audio employs metric learning techniques to train speaker embeddings that are optimized for a predefined distance metric, such as cosine similarity. This approach eliminates the need for additional steps like Probabilistic Linear Discriminant Analysis (PLDA), simplifying the clustering process.
3. End-to-End Training and Pipeline Integration:
pyannote.audio supports end-to-end training, allowing us to directly learn from the raw audio waveform without the need for manual feature extraction. This streamlines the workflow and enables the models to capture more nuanced acoustic information.
The pyannote.audio pipeline module integrates these modules into a cohesive framework, optimizing hyperparameters and jointly training the entire speaker segmentation pipeline. This approach, as opposed to independently training and combining individual modules, often leads to improved performance and more accurate speaker segmentation results.
4.3.2 Audio Description
In the Audio Description module, we leverage the powerful Whisper model to extract specific speech content from the audio segments. Whisper, an open-source speech recognition system developed by OpenAI, demonstrates exceptional performance across a wide range of languages and dialects. It excels in accurately transcribing speech, even in challenging acoustic environments. By employing end-to-end unsupervised training, Whisper efficiently learns from raw audio data, eliminating the need for manual annotations and significantly reducing training costs. Whisper’s versatility extends beyond speech recognition, as it can also be utilized for speech translation into multiple languages, enabling cross-lingual communication. Additionally, its ability to search for specific audio content based on keywords or phrases enhances its utility in information retrieval tasks. The modular design of Whisper, along with its availability in various sizes, allows for flexible deployment options based on specific needs and resource constraints. The open-source nature of Whisper further promotes collaboration and innovation within the speech recognition community, fostering the development of new applications and advancements in the field.
4.3.3 Emotion Analysis
@@
Emotion2Vec is an innovative deep learning based emotion recognition tool that aims to encapsulate the essence of emotional vocabulary in a rich high-dimensional vector space. This method not only captures the subtle relationships between words, but also assigns them the emotional burden of human expression.
The journey of Emotion2Vec begins with careful planning of the training dataset, where each word is carefully paired with its emotional essence - whether it’s discrete categories of happiness or anger, or subtle levels of intensity and arousal. From the perspective of models such as Word2Vec or GloVe, these words take on a new form, transformed into vectors that preserve their semantic structure while encoding their emotional meanings. As each emotional word gains its unique vector representation, a series of emotional vectors unfold, revealing the closeness of "happiness" to "joy" and its distance from "sadness". This mapping can provide a profound understanding of how emotions are integrated into language structures.
In the field of sentiment analysis, Emotion2Vec shines brightly, able to interpret emotions in various forms of data - whether it’s written text, spoken language, or visual images. It accurately classifies emotions, estimates their intensity, and identifies the most subtle similarities in emotions. The advantage of Emotion2Vec lies in its ability to grasp subtle semantic differences, express emotions clearly, and easily adapt to new emotional landscapes. Its applications cover the analysis of social media conversations, the classification of emotions in literary works, and the comparison of emotions between different datasets.
Ultimately, Emotion2Vec demonstrated the power of deep learning in the field of emotion recognition, providing emotional storytelling through vector language.The model in this article is derived from the Emotion2Vec related model trained by FunASR.
4.3.4 Word Speed
@@
3 This part is relatively simple, just calculate the number of words spoken by the speaker in a unit of time to obtain the speaking speed, using text descriptions such as "fast" and "slow"
@@ 5.1 Quantitative Analysis
Traditional NLP metrics such as BLEU (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), and CIDEr (Vedantam et al., 2015) evaluate generated text by comparing it to reference ground truth descriptions. While effective for assessing surface-level language quality, these metrics fail to capture the richness of knowledge beyond the visual domain, such as audio-related information, character details, or contextual nuances.
@@
Language Fluency: To assess language fluency, we utilize a keyword-triplet method that evaluates fluency at three levels of granularity: (1) within a single sentence, (2) between different sentences, and (3) the repetition of key-knowledge triplets relative to the overall story. Repetition of triplets within the story is considered a sign of unnatural phrasing and redundancy, which negatively impacts fluency. A higher fluency score is assigned when the text flows smoothly, with minimal repetition and a coherent structure.
Key-knowledge Relevance: For evaluating the relevance of key knowledge in the generated story, we define a knowledge base that includes both visual and audio information. In contrast to traditional models that focus solely on visual knowledge, our approach incorporates additional audio-related knowledge, such as ASR (Automatic Speech Recognition), environmental sound events, and other audio features. This expanded knowledge base allows us to assess the relevance and diversity of knowledge within the story. Specifically, we use two evaluation criteria: - Information Similarity (InfoSim): Measures the alignment between the knowledge points in the story and the knowledge repository. A higher similarity score indicates that the generated story effectively incorporates relevant knowledge. - Information Diversity (InfoDiverse): Evaluates the breadth of the knowledge used in the story. A higher diversity score indicates that the story incorporates a wide range of knowledge points, avoiding over-reliance on a small subset of information.
@@ 6 Conclusion
Acknowledgments
Appendix A Example Appendix
This is an appendix.
Model Fluency Info_sim Info_diverse Rouge_L BLEU-1 BLEU-2 BLEU-3 BLEU-4
Internvideo2 61.98 38.00 11.24 0.257 0.295 0.108 0.0265 0.0154
VideoLlava 15.94 38.18 13.64 0.376 0.275 0.113 0.0334 0.00474
M2s(GPT4o) 10.24 30.10 59.11 0.175 0.148 0.0556 0.0128 1.25e-6
M2s(llama) - - -
M2s(Flan-T5) - - -
Table 1: Performance of various models on Quantitative metrics
Model Environment Language Emotion Character SRCI SACOR_T Overall GPT Human
Internvideo2 1.0 1.0 0.0 0.6 0.2 0.600 0.56 -
VideoLlava 1.1 1.2 0.0 0.9 0.9 0.559 0.8 -
M2s(GPT4o) 2.7 2.3 1.0 1.7 1.7 0.735 1.9 -
M2s(llama) - - - - - -
M2s(Flan-T5) - - - - - -
Table 2: Performance of various models on Qualitative metrics
Model Fluency Info_sim Info_diverse Rouge_L BLEU-1 BLEU-2 BLEU-3 BLEU-4
M2s(without emo&speed) 7.31 29.87 52.26 0.196 0.166 0.0651 0.0137 1.24e-6
M2s(without ASR&emo&speed) 14.58 29.43 37.12 0.199 0.140 0.0626 0.0217 0.00365
M2s(GPT4o) 10.24 28.63 83.30 0.175 0.148 0.0556 0.0128 1.25e-6
Table 3: Performance of Ablation Study

### v2

@@ 4.2 Audio Feature Information Generation
Speaker Embedding for speaker recognition The core of our speaker recognition system lies in the Speaker Embedding module, powered by pyannote.audioet al. (2019). This module leverages deep learning models to extract unique and discriminative features from each speaker’s voice. These speaker embeddings serve as compact representations that capture the essential aspects of a speaker’s identity, facilitating accurate recognition and matching.
@@
Emotion Analysis Emotion2Vecet al. (2023c) is an innovative deep learning based emotion recognition tool that aims to encapsulate the essence of emotional vocabulary in a rich high-dimensional vector space. This method not only captures the subtle relationships between words, but also assigns them the emotional burden of human expression. Emotion2Vec demonstrated the power of deep learning in the field of emotion recognition, providing emotional storytelling through vector language. The model in this article is derived from the Emotion2Vec related model trained by FunASR.
@@
Word Speed This part is relatively simple, just calculate the number of words spoken by the speaker in a unit of time to obtain the speaking speed, using text descriptions such as "fast" and "slow".
4.3 LLMs selection and LoRA Finetuning
Different LLMs have varying abilities to integrate information and respond differently to instructions. Moreover, as a proprietary task, generic LLM may not necessarily have good results, so techniques such as LoRA may be needed to fine tune it. After extracting video and audio information, we can fine tune LoRAet al. (2021) for different LLMs to complete the final stage of generation and obtain stories.
Note that the data here can be real video audio caption story pairs generated through our pipeline. As introduced in the previous section, GPT can also be used to generate virtual data in the corresponding format to enhance data quality.
@@ 5.1 Quantitative Analysis
Traditional NLP metrics such as BLEUPapineni et al. (2002), ROUGELin (2004) evaluate generated text by comparing it to reference ground truth descriptions. While effective for assessing surface-level language quality, these metrics fail to capture the richness of knowledge beyond the visual domain, such as audio-related information, character details, or contextual nuances.
@@
Language Fluency: Fluency is used to evaluate the coherence and sentence repetition rate of generated explanatory texts. To assess language fluency, we utilize a keyword-triplet method that evaluates fluency at three levels of granularity: (1) within a single sentence, (2) between different sentences, and (3) the repetition of key-knowledge triplets relative to the overall story. Repetition of triplets within the story is considered a sign of unnatural phrasing and redundancy, which negatively impacts fluency. A higher fluency score is assigned when the text flows smoothly, with minimal repetition and a coherent structure. We have made modifications to the formulas in the original paperYang et al. (2024).Fluency and repetition are opposite. In the chart, we use repetition to represent the fluency of language, with lower repetition indicating smoother language. Then repetition can be represented by Intra-story Repetition (ISR): $\text{ISR}(S)=\frac{1}{n(n-1)}\sum_{i\neq j}\frac{\left|W(s_{i})\cap W(s_{j})\right|}{\left|W(s_{i})\right|}$ where $n$ is the number of explanatory texts, and $W(s_{i})$ represents the vocabulary set of explanatory text $s_{i}$. The formula calculates the repetition rate of vocabulary within the explanatory texts, which should be minimized to ensure diversity and coherence in the text.
Key-knowledge Relevance: For evaluating the relevance of key knowledge in the generated story, we define a knowledge base that includes both visual and audio information. In contrast to traditional models that focus solely on visual knowledge, our approach incorporates additional audio-related knowledge, such as ASR (Automatic Speech Recognition), environmental sound events, and other audio features. This expanded knowledge base allows us to assess the relevance and diversity of knowledge within the story. Specifically, we use two evaluation criteria: - Information Similarity (InfoSim)Yang et al. (2024): Measures the alignment between the knowledge points in the story and the knowledge repository. A higher similarity score indicates that the generated story effectively incorporates relevant knowledge. - Information Diversity (InfoDiverse)Yang et al. (2024): Evaluates the breadth of the knowledge used in the story. A higher diversity score indicates that the story incorporates a wide range of knowledge points, avoiding over-reliance on a small subset of information. The formula here is quite different from the original text. We have changed the normalization for time to the normalization for the number of keywords. That is to say, InfoDiverse calculates the proportion of keywords in the generated data that match the keywords in the reference data, ranging from 0 to 1.
InfoSim $=\frac{1}{2|s_{i}|}\sum_{s_{i}}\left(\max_{k\in K}f_{k}^{T}f_{s_{i}}\right.$
$\quad\left.+\frac{1}{|W(s_{i})|}\sum_{w\in W(s_{i})}\max_{k\in K}f_{k}^{T}f_{w}\right)$ (1)
InfoDiverse $=\frac{1}{|K|}\left|\bigcup_{s_{i}}\{k_{t}\in K\mid\max_{w\in W(s_{i})\cup S_{i}}\right.$
$\quad\left.f_{k_{t}}^{T}f_{w}>0.9\}\right|$ (2)
where $W(s_{i})$ represents all words in sentence $s_{i}$. $f_{s_{i}}$, $f_{k}$, and $f_{w}$ refer to the normalized embeddings of sentence $s_{i}$, knowledge point $k$, and segmented word $w$, respectively.
@@ 6 Conclusion
In this paper, we introduced MSBench, a comprehensive benchmark designed to evaluate the multi-modal story generation capabilities of MLLMs. We also proposed a robust MLLM pipeline, M2S, which outperforms leading models on the MSBench benchmark. Our extensive analyses provide valuable insights into the design of MLLMs for multi-modal story generation, particularly in scenarios enriched with additional information. Despite these advancements, our current approach has some limitations. We aim to address these in future work to enhance the evaluation framework and further refine the performance of MLLMs in complex, information-rich environments.
Appendix A Appendix Tables
Model ISR Info_sim Info_diverse SACOR Rouge-1 Rouge-2 Rouge_L BLEU-4
Internvideo2 15.1
VideoLlava 42.7
Videochat2 8.1 38.0 55.0
Internvideo2 + GPT4o 4.0 26.5 18.6 59.4 39.6 10.2 18.6 2.7
VideoLlava + GPT4o 5.6 25.0 20.7 61.1 40.9 10.5 19.0 3.2
Videochat2 + GPT4o 4.5 27.2 18.3 61.2 40.1 10.1 18.8 2.9
Internvideo2 + A-GPT4o 5.6 25.5 17.6 67.0
VideoLlava + A-GPT4o 6.0 25.2 21.9 65.3
Videochat2 + A-GPT4o 12.4 27.3 53.6 52.8
Table 1: Performance of various models with GPT4o as LLM on Quantitative metrics
Model Environment Character Emotion Language SRCI Overall GPT Human
Internvideo2 2.494 1.525 0.608 1.825 1.421 1.575
VideoLlava 2.170 1.155 0.213 1.539 1.118 1.239
Videochat2 3.086 1.878 0.921 2.243 1.809 1.987
Internvideo2 + GPT4o 3.766 2.675 1.792 3.555 2.717 2.901
VideoLlava + GPT4o 3.592 2.586 1.667 3.438 2.594 2.775
Videochat2 + GPT4o 3.873 2.651 1.738 3.550 2.727 2.908
Internvideo2 + A-GPT4o 3.581 2.574 1.659 3.433 2.585 2.767
VideoLlava + A-GPT4o 3.427 2.491 1.607 3.195 2.485 2.641
Videochat2 + A-GPT4o 3.492 2.568 1.673 3.090 2.625 2.689
Table 2: Performance of various models with GPT4o as LLM on Qualitative metrics
Model ISR Info_sim Info_diverse SACOR Rouge-1 Rouge-2 Rouge_L BLEU-4
GPT4o 12.4 27.3 53.6 52.8 - - - -
GPT3.5 13.5 26.3 25.5 54.0 38.2 8.6 17.6 1.9
Doubao 5.4 27.7 16.0 56.8 32.6 7.0 15.7 0.8
QWen 3.4 33.2 46.3 60.1 46.2 17.8 22.3 5.8
QWen-7B-instruct 9.8 35.3 50.8 58.3 21.4 34.8 16.5 27.3
mistral-instruct-7B-v1.0 8.1 8.3 7.5 3.5 4.7 1.8 6.3
llama3-3B 22.1 - 7.9 16.6 8.6 21. 24.6 8.2
lora_QWen-7B-instruct 7.6 33.1 50.8 58.2 52.4 21.7 25.1 9.0
lora_mistral-instruct-7B-v1.0 7.3 33.0 50.1 58.0 52.2 21.7 25.0 8.9
lora_llama3-3B 8.5 33.0 52.4 58.8 52.2 21.2 24.6 8.7
Table 3: Performance of various LLMs on Quantitative metrics
Model Environment Character Emotion Language SRCI Overall GPT Human
GPT4o 3.492 2.558 1.673 3.090 2.625 2.689 -
GPT3.5 2.831 1.809 1.077 2.299 1.815 1.966 -
Doubao 2.998 1.906 1.025 2.233 1.878 2.008 -
QWen 3.491 2.356 1.567 2.994 2.376 2.557 -
QWen-7B-instruct 2.826 1.884 1.125 2.224 1.805 1.973 -
mistral-instruct-7B-v1.0 0.360 0.229 0.124 0.297 0.213 0.245 -
llama3-3B 0.294 0.187 0.084 0.261 0.134 0.192 -
lora_QWen-7B-instruct 3.606 2.468 1.611 3.170 2.504 2.672 -
lora_mistral-instruct-7B-v1.0 3.623 2.494 1.625 3.159 2.527 2.686 -
lora_llama3-3B 3.560 2.431 1.577 3.106 2.457 2.626 -
Table 4: Performance of various LLMs on Qualitative metrics